PLEASE help me to find tables of value;
y=x^2+3x-3

Answer by uma(370)

Given y = x^2 + 3x - 3
To find the table of values, we subtitute different values for x and find the corresponding values for y.
(i) when x = 0
Plugging in this in the above we have..
y = 0 + 0 - 3
Therefore (0, -3) is an ordered pair
(ii) When x = 1,
y = 1^2 + 3(1) - 3
= 1 + 3 - 3
= 1
So (1,1) is another pair.
Thus for different vslues of x, we get different values for y.
Good Luck!!!